Hill Corporation issued $900,000 of 9% bonds at 99 on January 2, 2014. Interest is paid semiannually on June 30 and December 31. The bonds had a 10-year life from the date of issue, and the company uses the straight-line method of amortization. On March 31, 2017, Hill recalls the bonds at the call price of 107 plus accrued interest.

Prepare the journal entries to record the reacquisition (recall) of Hill’s bonds.
Answer #1

Discount on bonds payable = 900000*1/100 = 9000

Discount amortization = 9000*39/120 = 2925

Unamortized discount = 9000-2925 = 6075

Journal entries

Date account and explanation debit credit
Mar 31,2017 Interest expense 20475
Discount on bonds payable (9000*3/120) 225
Cash (900000*9%*3/12) 20250
(To record interest)
Mar 31,2017 Bonds payable 900000
Loss on redemption of bonds 69075
Discount on bonds payable 6075
Cash (900000*1.07) 963000
(To record redemption)